OpenStudy (anonymous):

Find the surface area for the right prism. Base = 10m Height of the Base = 12m Height = 12m

OpenStudy (anonymous):

Here is a picture of the problem.

OpenStudy (shamim):

u hv 3 rectangle. right?

OpenStudy (shamim):

u hv 2 triangle

OpenStudy (anonymous):

Yes, am I supposed to split them in half?

OpenStudy (shamim):

lenght of ur rectangle is 12m nd width is 10 m

OpenStudy (valpey):

The surface area will be the sum of the areas of each side. I think you might need to assume the ends are isosceles in which case you have two 5-12-13 triangles back-to-back.

OpenStudy (shamim):

now area of 1 rectangle is 12m*10m

OpenStudy (anonymous):

I calculated 120m.

OpenStudy (anonymous):

At first, I thought I was supposed to split the triangle in half, so that each one has a base of 5 (10/2). But there's still one side missing, and I don't know how to find it.

OpenStudy (shamim):

120m^2

OpenStudy (anonymous):

120m^2=14400.

OpenStudy (shamim):

so the total area of 3 rectangle is 3*120=360m^2

OpenStudy (shamim):

right?

OpenStudy (anonymous):

Yes.

OpenStudy (shamim):

now we hv to find out the area of tringle

OpenStudy (shamim):

area of a triangle=1/2*base*height

OpenStudy (anonymous):

60?

OpenStudy (shamim):

ares of a triangle=1/2*10*12=60m^2

OpenStudy (shamim):

ya u r right

OpenStudy (shamim):

now tell me the total area of ur two triangle

OpenStudy (anonymous):

120m?

OpenStudy (shamim):

now just add total area of ur 2 triangle with total area of ur rectangle

OpenStudy (shamim):

no

OpenStudy (anonymous):

480m^2?

OpenStudy (shamim):

ya u r successful

OpenStudy (anonymous):

Thank you!

OpenStudy (shamim):

u r most welcome

OpenStudy (valpey):

Uh, no. The three rectangles are (12x10, 12x13, and 12x13) The two triangles do add up to 120 m^2, though, so that is right.

OpenStudy (valpey):

You could use the Pythagorean Theorum and the isoceles assumption to arrive at the other edge lengths being equal to 13.